# Differentiable and accelerated spherical harmonic and Wigner transforms

### v1

@@ 3.1 Explicit Wigner representations
A number of explicit representations of the $d$-functions can be considered, such as the differential representation (Varshalovich et al., 1989, Sec. 4.3.2, p. 77, Eq. 7)
@@
or the expression in terms of Jacobi polynomials $P_{s}^{\mu,\nu}(\cdot)$ given by (Varshalovich et al., 1989, Sec. 4.3.4, p. 78, Eq. 13–15)
@@ 3.2 Requirements for parallel computation and automatic differentiation
A number of recursions, while stable, recurse over both harmonic degree $\ell$ and order $m$ (e.g. Risbo, 1996; Trapani and Navaza, 2006). This limits the degree of parallelisation that can be achieved. Three-term recurrences typically recurse over a single index only, which is well-suited for high levels of parallelisation, although they can be unstable with errors compounding and resulting in highly inaccurate calculation Gautschi (1967). Nevertheless, careful attention can be given to ensure stability can be achieved (e.g. Prézeau and Reinecke, 2010; Reinecke and Seljebotn, 2013).
@@ 3.3 Recursion
Our algorithm is built upon the three-term Wigner $d$-function recursion (Varshalovich et al., 1989, Sec. 4.8.3, p. 93, Eq. 17)
@@
This recursion iterates across order $m$ independently of $\ell$, $n$, and $\beta$, thereby supporting an algorithmic structure that can be highly parallelised. While computing this recursion for increasing $m$ is unstable, it can be computed stably for decreasing $m$ (as discussed in detail in Prézeau and Reinecke (2010), although they consider recursion over $n$). One may straightforwardly rearrange Eq. 24 to recover
@@
The recursion needs to be initialised at $d^{\ell}_{\ell n}(\beta)$ so that we can recurse down in $m$, starting from $m=\ell$. We therefore initialise the algorithm using the relation
@@
which follows by the expression for the Wigner $d$-functions in terms of Jacobi polynomials of Eq. 22. To sketch how one reaches this result, consider the case in which the spin number $n=\ell$, the Jacobi polynomial term then reduces to unity, and the symmetry relation $d^{\ell}_{mn}(\beta)=(-1)^{n-m}d^{\ell}_{nm}(\beta)$ (Varshalovich et al., 1989, Sec. 4.4, p. 79, Eq. 1) can then be applied to reorder the $m$ and $n$ indices, following which the trigonometric expressions and prefactor follow straightforwardly. We calculate these terms in log-space to mitigate numerical issues inherent in evaluating large factorials. Also note that $d^{\,\ell}_{\ell+1,n}(\beta)$ is trivially zero as $d^{\,\ell}_{mn}(\beta)=0$ for all $m>\ell$.
While some spherical samplings avoid samples at the north and south poles, others do have samples on the poles. Since we aim to support a number of different spherical samplings (as discussed in Sec. 2.5), including those with samples defined on the poles, we must be able to evaluate Wigner $d$-functions on the poles. The initialisation and recursion presented above are not suitable on the poles ($\beta=\{0,\pi\}$) since the recursion coefficient $\lambda_{m}$ is undefined due to the $\sin\beta$ term in the denominator. Fortunately, simple forms for these elements of the Wigner $d$-functions are available, which we implement explicitly. Specifically, at the north pole (Varshalovich et al., 1989, Sec. 4.16, p. 112, Eq. 1)
@@
which follows from Varshalovich et al. (1989, Sec. 4.4, p. 79, Eq. 1). Our recursive algorithm to compute the Wigner $d$-functions in a highly parallelised and stable manner, and its initialisation, is outlined graphically in Fig. 2 for clarity.
{tikzpicture}
@@
Additionally, note that the Wigner $d$-functions exhibit an extreme dynamic range. For example, the magnitude of these coefficients $d^{\ell}_{mn}(\beta)$ often ranges in excess of many hundreds of orders of magnitude. Hence, not only must one recurse along stable paths but one must also develop protocols to avoid under- and over-flowing, which can be equally catastrophic. A common approach to address this issue is to implement a renormalisation scheme, which conditionally compresses the dynamic range during the recursion to an acceptable interval. However, while this may solve the dynamic range problem in theory, algorithms with this conditional structure do not permit efficient automatic differentiation. In fact, due to branching trees, such an algorithm can become extremely expensive to differentiate. Alternative methods exist, such as continued fractions (e.g. Gautschi, 1967), however, as we discuss below in Sec. 3.4, with some careful software engineering it is possible to develop non-conditional algorithms for renormalisation that are not only feasible, but exhibit preferable complexity scaling.
@@ 4 Fast Fourier transforms on the sphere and rotation group
In much the same way that the standard fast Fourier transform (FFT; Cooley and Tukey, 1965) is reliant on uniform sampling, fast generalised Fourier transforms on the sphere and rotation group are reliant on isolatitudinal sampling. Consequently, we focus our attention solely to isolatitudinal sampling (as discussed in Sec. 2.4), which includes the majority of commonly adopted spherical sampling approaches (as discussed in Sec. 2.5). The key advantage of isolatitudinal samplings is that a separation of variables may be performed to reduce computational complexity (e.g. Górski et al., 2005; McEwen and Wiaux, 2011; McEwen et al., 2015, 2013).
@@ 4.2.1 Driscoll & Healy (DH) sampling theorem
The DH (Driscoll & Healy; Driscoll and Healy, 1994) sampling theorem adopts equiangular sampling position on the sphere defined by
@@
resulting in $\sim 4L^{2}$ samples on the sphere. These sample locations afford a sampling theorem with explicit quadrature weights for spherical transforms given by Driscoll and Healy (1994); McEwen et al. (2011)
@@
Spherical harmonic transforms can thus be computed straightforwardly for the DH sampling theorem, following the algorithmic structure outlined above in Sec. 4.1.
@@ 4.2.2 McEwen & Wiaux (MW) sampling theorem
The MW (McEwen & Wiaux; McEwen and Wiaux, 2011) sampling theorem provides an alternative equiangular sampling theorem requiring only $\sim 2L^{2}$ samples on the sphere, providing a reduction in the Nyquist rate on the sphere by a factor of two compared to the DH sampling theorem.
There are two variants of the MW sampling scheme: the original scheme which provides the minimal Nyquist sampling on the sphere McEwen and Wiaux (2011); and a variant which slightly over samples (but yet still with $\sim 2L^{2}$ samples) to yield sample locations with antipodal symmetry that is useful for many practical applications Daducci et al. (2011); Ocampo et al. (2023), denoted MWSS (MW with symmetric sampling). The MW sample positions are given by
@@
Quadrature weights have been presented for both the MW McEwen and Wiaux (2011) and MWSS Ocampo et al. (2023) cases, which we unify into a single expression. The MW sampling theorem is based on an extension of the sphere to the torus through careful period extensions of the function of interest to the $\theta$ domain $[0,2\pi)$; recall that on the sphere $\theta\in[0,\pi]$. Consider the bandlimited representation of the function defined by $\sin\theta$ on $[0,\pi]$ and zero on $(\pi,2\pi)$, given by
@@
Signals sampled following either of the MW sampling approaches can first be extended to the torus by a reflection in $\theta$, combined with the introduction on the extended domain of a shift of $\pi$ in $\phi$ and a $(-1)^{s}$ sign shift McEwen and Wiaux (2011). While this extension can be computed in harmonic space for either MW or MWSS sampling theorems, it can also be straightforwardly computed in the spatial domain for MWSS due to the symmetric nature of the sample positions. Once the extended function on the torus is computed it can be upsampling simply by zero padding in the Fourier domain and then truncated back to the spherical domain $\theta\in[0,\pi]$. Following this approach the algorithmic structure of the spherical harmonic transform algorithms presented in Sec. 4.1 can be followed directly.
@@ 4.2.3 Gauss-Legendre (GL) sampling theorem
An alternative sampling theorem on the sphere can be built on Gauss-Legendre quadrature, where sample positions in $\theta$ are defined by the roots of the Legendre polynomials of order $L$ (e.g. Skukowsky, 1986; McEwen and Wiaux, 2011). Samples positions in $\phi$ may be defined by
@@
While equiangular sampling theorems are usually preferred for practical purposes due to the regular nature of sample positions in $\theta$, the GL sampling theorem on the sphere recovers the same asymptotic number of samples on the sphere as the MW sampling theorem ($\sim 2L^{2}$), although in practice slightly more samples are required by the GL sampling compared to MW. Explicit quadrature weights for the computation of spherical transforms are given by Press et al. (1992)
@@
where above the prime denotes the derivative of the Legendre polynomial. Spherical harmonic transforms can thus by computed straightforwardly for the GL sampling theorem, following the algorithmic structure outlined above in Sec. 4.1.
@@ 4.2.4 HEALPix sampling scheme
HEALPix sampling is designed to provide pixels of equal area, which can be a very useful practical property. We refer the reader to Górski et al. (2005) for full details regarding HEALPix, including the precise definition of sample positions, and focus here on the implications for computing spherical harmonic transforms. Since pixels have equal area, approximate quadrature weights are simply given by $q(\theta,\phi)=4\pi/N_{\text{pix}}$ where $N_{\text{pix}}$ denotes the total number of samples over the sphere. Before proceeding, note that for HEALPix it is customary to compute transforms to a maximum degree denoted by $\ell_{\text{max}}$, which is related to our definition of bandlimit $L$ by $\ell_{\text{max}}=L-1$.
@@
Fig. 3: Overview of the additional steps required for HEALPix sampling Górski et al. (2005) when computing Fourier transforms over $\phi$. For HEALPix one must: (i) compute the Fourier transform each ring at varying resolution, determined by $n_{\phi}$ per ring; then (ii) generate synthetic orders $m^{\prime}$ by extending periodically available orders $|\,m\,|\leq|\,m^{\prime}\,|$. For the inverse spherical harmonic transform the order is reversed and step (ii) is simply replaced by aliasing higher orders $m^{\prime}$ $|\,m^{\prime}\,|>|\,m\,|$ into lower orders $m$. This is in fact precisely why HEALPix provides only approximate spherical harmonic transforms. For equirectangular sampling schemes this entire step can be computed simply by an FFT for each $\phi$ for all rings in a single line of code.
@@ 4.3 Full precomputation
We now consider how best to evaluate the summations involving the Wigner $d$-functions presented in Eq. 30 and Eq. 33. For moderate bandlimits ($L\lesssim 2048$) the Wigner $d$-functions may be precomputed, yielding highly efficient harmonic transforms, although at a cost of increased memory usage. One may precompute and store all required elements $d^{\ell}_{m,-s}(\theta)$, the weighted sum of which can then be implemented extremely straightforwardly. Explicitly, for a forward spherical harmonic transform, given the precomputation and storage of the Wigner $d$-functions $\boldmath{D}=d^{\ell}_{m,-s}(\theta)$ for all $\ell,m,\theta$ at a fixed spin $s$, the transform is simply given by a series of 1D FFTs over $\phi$ for each $\theta$, followed by multiplication by the quadrature weights, followed by multiplication by matrix $\boldmath{D}$, followed by an $\ell$-dependent scaling. The core implementation of the spherical harmonic transform therefore collapses effectively to a single line of code jnp.einsum(‘‘tlm, tm -> lm’’, $\boldmath{D}$, jnp.fft(f)). The inverse transform follows analogously, with minor changes (different ordering, forward 1D FFTs replaced by inverse FFTs, and different scalings). While this precompute approach is highly computationally efficient, it requires $\mathcal{O}(L^{3})$ memory, which is not feasible at high bandlimits $L$. We next consider an approach that avoids fully precomputing Wigner $d$-functions, eliminating $\mathcal{O}(L^{3})$ memory requirements, and thus is scalable to high degrees $L$. In this approach we do consider an optional precomputation of some other internal terms, although in a manner that limits memory requirements to $\mathcal{O}(L^{2})$.
@@ 4.4 Interleaving recursion and computation for on-the-fly transforms
{tikzpicture}
@@ 4.5 Wigner transforms
To conclude this section we consider the computation of the forward Wigner transform of bandlimited signals $f\in\mathcal{B}_{L}({\text{SO}(3)})$ denoted by $\mathcal{W}:\mathcal{B}_{L}(\text{SO}(3))\rightarrow\mathbb{C}^{(4L^{3}-L)/3}$ and the inverse denoted by $\mathcal{W}^{-1}:\mathbb{C}^{(4L^{3}-L)/3}\rightarrow\mathcal{B}_{L}(\text{SO}(3))$. As discussed in Sec. 2.3 (McEwen et al., 2015) by relating the Wigner $D$-functions to the spin spherical harmonic basis functions, one can compute the Wigner transform through a series of spin spherical harmonic transforms and Fourier transforms (see Eq. 15–Eq. 18). We can therefore make use of the approach previously discussed to compute forward and inverse spherical harmonic transforms for arbitrary spin, making the association $\alpha\leftrightarrow\phi$ and $\beta\leftrightarrow\theta$, and simply include an additional 1D FFT with respect to $\gamma$.

### v2

@@ 3.1 Explicit Wigner representations
A number of explicit representations of the $d$-functions can be considered, such as the differential representation [48, Sec. 4.3.2, p. 77, Eq. 7]
@@
or the expression in terms of Jacobi polynomials $P_{s}^{\mu,\nu}(\cdot)$ given by [48, Sec. 4.3.4, p. 78, Eq. 13–15]
@@ 3.2 Requirements for parallel computation and automatic differentiation
A number of recursions, while stable, recurse over both harmonic degree $\ell$ and order $m$ [e.g. 57, 58]. This limits the degree of parallelisation that can be achieved. Three-term recurrences typically recurse over a single index only, which is well-suited for high levels of parallelisation, although they can be unstable with errors compounding and resulting in highly inaccurate calculation [59]. Nevertheless, careful attention can be given to ensure stability can be achieved; for example, often recursing in some directions can be highly unstable, while other directions may be stable [e.g. 59, 60, 61].
@@ 3.3 Recursion
Our algorithm is built upon the three-term Wigner $d$-function recursion [48, Sec. 4.8.3, p. 93, Eq. 17]
@@
This recursion iterates across order $m$ independently of $\ell$, $n$, and $\beta$, thereby supporting an algorithmic structure that can be highly parallelised. While computing this recursion for increasing $m$ is unstable, it can be computed stably for decreasing $m$ (as discussed in detail in Prézeau and Reinecke [60], although they consider recursion over $n$). One may straightforwardly rearrange Eq. 24 to recover
@@
The recursion needs to be initialised at $d^{\ell}_{\ell n}(\beta)$ so that we can recurse down in $m$, starting from $m=\ell$. For this particular relation, recursing along decreasing harmonic order $m$ is stable, whereas the increasing $m$ direction causes numerical errors to compound. We therefore initialise the algorithm using the relation
@@
which follows by the expression for the Wigner $d$-functions in terms of Jacobi polynomials of Eq. 22. To sketch how one reaches this result, consider the case in which the spin number $n=\ell$, the Jacobi polynomial term then reduces to unity, and the symmetry relation $d^{\ell}_{mn}(\beta)=(-1)^{n-m}d^{\ell}_{nm}(\beta)$ [48, Sec. 4.4, p. 79, Eq. 1] can then be applied to reorder the $m$ and $n$ indices, following which the trigonometric expressions and prefactor follow straightforwardly. We calculate these terms in log-space to mitigate numerical issues inherent in evaluating large factorials. Also note that $d^{\,\ell}_{\ell+1,n}(\beta)$ is trivially zero as $d^{\,\ell}_{mn}(\beta)=0$ for all $m>\ell$.
While some spherical samplings avoid samples at the north and south poles, others do have samples on the poles. Since we aim to support a number of different spherical samplings (as discussed in Sec. 2.5), including those with samples defined on the poles, we must be able to evaluate Wigner $d$-functions on the poles. The initialisation and recursion presented above are not suitable on the poles ($\beta=\{0,\pi\}$) since the recursion coefficient $\lambda_{m}$ is undefined due to the $\sin\beta$ term in the denominator. Fortunately, simple forms for these elements of the Wigner $d$-functions are available, which we implement explicitly. Specifically, at the north pole [48, Sec. 4.16, p. 112, Eq. 1]
@@
which follows from Varshalovich et al. [48, Sec. 4.4, p. 79, Eq. 1]. Our recursive algorithm to compute the Wigner $d$-functions in a highly parallelised and stable manner, and its initialisation, is outlined graphically in Fig. 2 for clarity.
Refer to caption Initialisation (see Eq. 27)
Refer to caption Recursion (see Eq. 25)
@@
Additionally, note that the Wigner $d$-functions exhibit an extreme dynamic range. For example, the magnitude of these coefficients $d^{\ell}_{mn}(\beta)$ often ranges in excess of many hundreds of orders of magnitude. Hence, not only must one recurse along stable paths but one must also develop protocols to avoid under- and over-flowing, which can be equally catastrophic. A common approach to address this issue is to implement a renormalisation scheme, which conditionally compresses the dynamic range during the recursion to an acceptable interval. However, while this may solve the dynamic range problem in theory, algorithms with this conditional structure do not permit efficient automatic differentiation. In fact, due to branching trees, such an algorithm can become extremely expensive to differentiate. Alternative methods exist, such as continued fractions [e.g. 59], however, as we discuss below in Sec. 3.4, with some careful software engineering it is possible to develop non-conditional algorithms for renormalisation that are not only feasible, but exhibit preferable complexity scaling.
@@ 4 Fast Fourier transforms on the sphere and rotation group
In much the same way that the standard fast Fourier transform [FFT; 62] is reliant on uniform sampling, fast generalised Fourier transforms on the sphere and rotation group are reliant on isolatitudinal sampling. Consequently, we focus our attention solely to isolatitudinal sampling (as discussed in Sec. 2.4), which includes the majority of commonly adopted spherical sampling approaches (as discussed in Sec. 2.5). The key advantage of isolatitudinal samplings is that a separation of variables may be performed to reduce computational complexity [e.g. 43, 41, 42, 50].
@@ 4.2.1 Driscoll & Healy (DH) sampling theorem
The DH [Driscoll & Healy; 40] sampling theorem adopts equiangular sampling position on the sphere defined by
@@
resulting in $\sim 4L^{2}$ samples on the sphere. These sample locations afford a sampling theorem with explicit quadrature weights for spherical transforms given by [40, 63]
@@
Note that the quadrature weights are constant with respect to $\phi$. Spherical harmonic transforms can thus be computed straightforwardly for the DH sampling theorem, following the algorithmic structure outlined above in Sec. 4.1.
@@ 4.2.2 McEwen & Wiaux (MW) sampling theorem
The MW [McEwen & Wiaux; 41] sampling theorem provides an alternative equiangular sampling theorem requiring only $\sim 2L^{2}$ samples on the sphere, providing a reduction in the Nyquist rate on the sphere by a factor of two compared to the DH sampling theorem.
There are two variants of the MW sampling scheme: the original scheme which provides the minimal Nyquist sampling on the sphere [41]; and a variant which slightly over samples (but yet still with $\sim 2L^{2}$ samples) to yield sample locations with antipodal symmetry that is useful for many practical applications [6, 64], denoted MWSS (MW with symmetric sampling). The MW sample positions are given by
@@
Quadrature weights have been presented for both the MW [41] and MWSS [64] cases, which we unify into a single expression. The MW sampling theorem is based on an extension of the sphere to the torus through careful period extensions of the function of interest to the $\theta$ domain $[0,2\pi)$; recall that on the sphere $\theta\in[0,\pi]$. Consider the bandlimited representation of the function defined by $\sin\theta$ on $[0,\pi]$ and zero on $(\pi,2\pi)$, given by
@@
Signals sampled following either of the MW sampling approaches can first be extended to the torus by a reflection in $\theta$, combined with the introduction on the extended domain of a shift of $\pi$ in $\phi$ and a $(-1)^{s}$ sign shift [41]. While this extension can be computed in harmonic space for either MW or MWSS sampling theorems, it can also be straightforwardly computed in the spatial domain for MWSS due to the symmetric nature of the sample positions. Once the extended function on the torus is computed it can be upsampling simply by zero padding in the Fourier domain and then truncated back to the spherical domain $\theta\in[0,\pi]$. Following this approach the algorithmic structure of the spherical harmonic transform algorithms presented in Sec. 4.1 can be followed directly.
@@ 4.2.3 Gauss-Legendre (GL) sampling theorem
An alternative sampling theorem on the sphere can be built on Gauss-Legendre quadrature, where sample positions in $\theta$ are defined by the roots of the Legendre polynomials of order $L$ [e.g. 54, 41]. Samples positions in $\phi$ may be defined by
@@
While equiangular sampling theorems are usually preferred for practical purposes due to the regular nature of sample positions in $\theta$, the GL sampling theorem on the sphere recovers the same asymptotic number of samples on the sphere as the MW sampling theorem ($\sim 2L^{2}$), although in practice slightly more samples are required by the GL sampling compared to MW. Explicit quadrature weights for the computation of spherical transforms are given by [65]
@@
where above the prime denotes the derivative of the Legendre polynomial.The quadrature weights are constant with respect to $\phi$. Spherical harmonic transforms can thus by computed straightforwardly for the GL sampling theorem, following the algorithmic structure outlined above in Sec. 4.1.
@@ 4.2.4 HEALPix sampling scheme
HEALPix sampling is designed to provide pixels of equal area, which can be a very useful practical property. We refer the reader to Górski et al. [43] for full details regarding HEALPix, including the precise definition of sample positions, and focus here on the implications for computing spherical harmonic transforms. Since pixels have equal area, approximate quadrature weights are simply given by $q(\theta,\phi)=4\pi/N_{\text{pix}}$ where $N_{\text{pix}}$ denotes the total number of samples over the sphere. The quadrature weights are constant with respect to both $\theta$ and $\phi$. Before proceeding, note that for HEALPix it is customary to compute transforms to a maximum degree denoted by $\ell_{\text{max}}$, which is related to our definition of bandlimit $L$ by $\ell_{\text{max}}=L-1$.
@@
Fig. 3: Overview of the additional steps required for HEALPix sampling [43] when computing Fourier transforms over $\phi$. For HEALPix one must: (i) compute the Fourier transform each ring at varying resolution, determined by $n_{\phi}$ per ring; then (ii) generate synthetic orders $m^{\prime}$ by extending periodically available orders $|\,m\,|\leq|\,m^{\prime}\,|$. For the inverse spherical harmonic transform the order is reversed and step (ii) is simply replaced by aliasing higher orders $m^{\prime}$ $|\,m^{\prime}\,|>|\,m\,|$ into lower orders $m$. This is in fact precisely why HEALPix provides only approximate spherical harmonic transforms. For equirectangular sampling schemes this entire step can be computed simply by an FFT for each $\phi$ for all rings in a single line of code.
@@ 4.3 Full precomputation
We now consider how best to evaluate the summations involving the Wigner $d$-functions presented in Eq. 30 and Eq. 33. For moderate bandlimits ($L\lesssim 2048$) the Wigner $d$-functions may be precomputed, yielding highly efficient harmonic transforms, although at a cost of increased memory usage. One may precompute and store all required elements $d^{\ell}_{m,-s}(\theta)$, the weighted sum of which can then be implemented extremely straightforwardly. Explicitly, for a forward spherical harmonic transform, given the precomputation and storage of the Wigner $d$-functions $\boldmath{D}=d^{\ell}_{m,-s}(\theta)$ for all $\ell,m,\theta$ at a fixed spin $s$, the transform is simply given by a series of 1D FFTs over $\phi$ for each $\theta$, followed by multiplication by the quadrature weights, followed by multiplication by matrix $\boldmath{D}$, followed by an $\ell$-dependent scaling. The core implementation of the spherical harmonic transform therefore collapses effectively to a single line of code jnp.einsum(‘tlm, tm -> lm’, $\boldmath{D}$, jnp.fft(f)). The inverse transform follows analogously, with minor changes (different ordering, forward 1D FFTs replaced by inverse FFTs, and different scalings). While this precompute approach is highly computationally efficient, it requires $\mathcal{O}(L^{3})$ memory, which is not feasible at high bandlimits $L$. We next consider an approach that avoids fully precomputing Wigner $d$-functions, eliminating $\mathcal{O}(L^{3})$ memory requirements, and thus is scalable to high degrees $L$. In this approach we do consider an optional precomputation of some other internal terms, although in a manner that limits memory requirements to $\mathcal{O}(L^{2})$.
@@ 4.4 Interleaving recursion and computation for on-the-fly transforms
Refer to caption $d^{\ell}_{mn}(\theta)$${}_{s}\hat{f}_{\ell m}$${}_{s}\tilde{f}_{m}(\theta)$${}_{s}f(\theta,\phi)$Eq. 32
[…]
Eq. 31
[…]
$\sum\limits_{\ell}$q. 33
[…]
$\times$
[…]
$\sum\limits_{\theta}$q. $\mathcal{F}$$\mathcal{F}^{-1}$
@@ 4.5 Wigner transforms
To conclude this section we consider the computation of the forward Wigner transform of bandlimited signals $f\in\mathcal{B}_{L}({\text{SO}(3)})$ denoted by $\mathcal{W}:\mathcal{B}_{L}(\text{SO}(3))\rightarrow\mathbb{C}^{(4L^{3}-L)/3}$ and the inverse denoted by $\mathcal{W}^{-1}:\mathbb{C}^{(4L^{3}-L)/3}\rightarrow\mathcal{B}_{L}(\text{SO}(3))$. As discussed in Sec. 2.3 [42] by relating the Wigner $D$-functions to the spin spherical harmonic basis functions, one can compute the Wigner transform through a series of spin spherical harmonic transforms and Fourier transforms (see Eq. 15–Eq. 18). We can therefore make use of the approach previously discussed to compute forward and inverse spherical harmonic transforms for arbitrary spin, making the association $\alpha\leftrightarrow\phi$ and $\beta\leftrightarrow\theta$, and simply include an additional 1D FFT with respect to $\gamma$.